OpenStudy (kabase1234):

Please help I'll fan and medal Find the first six terms of the sequence. a1 = -8, an = 5 • an-1

OpenStudy (kabase1234):

@mathmate

OpenStudy (kabase1234):

@marc_d

OpenStudy (mathmate):

The rule is an = 5 • an-1 which means a1=-8 a2=5(-8)=-40 a3=5^2(-8)=-200 a4=... And in general, an=5^(n-1)(-8) So there should be enough information for you to calculate a1 to a6.

OpenStudy (kabase1234):

I'm new at these and I don't know how to do them that much

OpenStudy (kabase1234):

Wait is the first term of the sequence -8?

OpenStudy (kabase1234):

Wait! Is it -8, -40, -200, -1000, -5000, -25,000

OpenStudy (kabase1234):

@mathmate

OpenStudy (mathmate):

correct! Well done! :)